The relative electrophoretic mobilities of a 30-kDa protein and a 92-kDa protein used as standards on a SDS-polyacrylamide gel are 0.80 and 0.41, respectively. What is the apparent mass of a protein having a mobility of 0.62 on this gel.
Answer #1

Sol :-

From equation of straight line :

Y = AX + B ................(1)

Here, X = protein mobility = 0.62 (Given)

Y = y-axis on which log(M) is taken , M = apparent mass of a protein = ?

X = x-axis on which Rf (mobility) is taken and

A and B are constants whose values have to calculate

Now,

When apparent mass of protein is 30, then mobility (Rf) is 0.80 , So, equation (1) becomes :

log 30 = 0.80 A + B

1.477 = 0.80 A + B .............................(2)

When apparent mass of protein is 92, then mobility (Rf) is 0.41 , So, equation (1) becomes :

log 92 = 0.41 A + B

1.964 = 0.41 A + B .............................(3)

On subtracting equation (2) from equation (3) :

1.964 - 1.477 = 0.41 A - 0.80 A + B - B

0.487 = - 0.39 A

A = - 0.487/0.39

A = - 1.25

Substitute the value of A in equation (2) :

1.477 = 0.80 (-1.25) + B

B = 1.477 +1

B = 2.48

Substitute the values of A, B and X in equation (1)

Y = log(M) = (-1.25)(0.62) + 2.48

log(M) = - 0.775 + 2.48

log(M) = 1.705

M = 101.705

M = 50.7 -KDa

Hence, apparent mass of a protein = M = 50.7-kD